Given f(x,y) = 2 ; 0 <X<y< 1 a. Prove that f(x,y) is a joint pdf b. Find the correlation coefficient of X and Y
Answer #1

In this for obtaining the correlation we first find marginal distribution function and then using marginal we can obtain the covariance and hence correlation can be obtained. Here the limit of x and y depend on one another .

frnigd is a pdf Xy We have flu,y) = 2 of us yaa. 9. If 5 sf (my) obedy Boty - Y=2 > 1 [may (1) de 25 adu dy = fa 2 dn sady =- 2x(4-5) = 2* (34) = {(t) = 1 tyly) ay = 1 & 2 yo y dy u(x) = E(X2) - É() 12 {(x²) = 52 m². fx (m) dhe = $.261-10) du = $50N(Y) = = {(7) - (E(Y) 22 -3)2 = 0.5-0.444 - 0.055. Also, {(xx) = go dt (xxy) fxx (M14) don day = 8 2 1h sry 12, dn dy = 5² 27